OpenStudy (anonymous):

find the derivative f(x)=(c^9-x^9)/(c^9+x^9), c is constant

geerky42 (geerky42):

What do you need help with?

OpenStudy (anonymous):

its calculus find the derivation which i know how to do but i have no clue what c is constant means

geerky42 (geerky42):

c is just a unknown number. treat it as number.

OpenStudy (campbell_st):

image it's 512 or 2^9 that's all it means... some arbitrary number

OpenStudy (anonymous):

if i just take the derivation of that equation though using the quotient formula it will equal out to 0 which is wrong

OpenStudy (freckles):

derivation?

OpenStudy (freckles):

derivative?

OpenStudy (anonymous):

yes derivative sorry

OpenStudy (freckles):

would you know how to find the derivative of say something like: \[f(x)=\frac{1^9-x^9}{1^9+x^9}=\frac{1-x^9}{1+x^9}\]

OpenStudy (anonymous):

yes i just use the quotient rule I just don't understand what c being constant means

OpenStudy (anonymous):

or i can use the chain rule to figure it out

OpenStudy (freckles):

like c doesn't change c doesn't vary it isn't a variable derivative of c is 0

OpenStudy (freckles):

just like the derivative of 1 is 0

OpenStudy (freckles):

1 is actually called a constant because it constantly remains 1

geerky42 (geerky42):

\(\dfrac{d}{dx}c = 0\)

OpenStudy (freckles):

sqrt(30) is a constant

OpenStudy (freckles):

i'm sure you heard of rules like constant multiple rule and constant rule if you are being asked to differentiate this

OpenStudy (freckles):

\[(cf)'=c(f')\] means if c is some number like 5 or 3 or -2 or -1/2 or sqrt(3) then you can factor it out and bring down side the derivative of the part that varies

OpenStudy (anonymous):

actually no my teacher is a ta thats not so great lol i teach myself everything

OpenStudy (freckles):

and that was the constant multiple rule i just mentioned

OpenStudy (freckles):

the constant rule is derivative of a constant being 0

OpenStudy (anonymous):

so basically i replace c with 0

OpenStudy (freckles):

\[f(x)=\frac{c^9-x^9}{c^9+x^9} \\ f'(x)=\frac{(c^9-x^9)'(c^9+x^9)-(c^9+x^9)'(c^9-x^9)}{(c^9+x^9)^2}\] so basically since c is a constant c^9 is a constant and derivative of c^9 would be 0

OpenStudy (anonymous):

and if i do that the derivative will still all equal to 0 which web assign is telling me is wrong because i originally thought that you would take c equal it to 0 but when web assign said it was wrong i just assume I didn't know what c being constant really meant

OpenStudy (freckles):

no the derivative of all of that is not equal to 0

OpenStudy (freckles):

what is derivative of (c^9-x^9)?

OpenStudy (anonymous):

9c^8-9x^8

OpenStudy (freckles):

\[\frac{d}{dx}(c^9-x^9)=\frac{d}{dx}c^9 -\frac{d}{dx} x^9\] what does this equal?

OpenStudy (freckles):

not exactly

OpenStudy (freckles):

again c is a constant

OpenStudy (freckles):

\[\frac{d}{dx}(c^9-x^9)=\frac{d}{dx}c^9 -\frac{d}{dx}x^9=0-9x^8 \]

OpenStudy (anonymous):

well if you are talking about it being constant then c would be 0 so it will be -9x^8

OpenStudy (freckles):

well c wouldn't be zero but the derivative of c is zero

OpenStudy (anonymous):

ok i get it now thanks

OpenStudy (freckles):

\[f(x)=\frac{c^9-x^9}{c^9+x^9} \\ f'(x)=\frac{(c^9-x^9)'(c^9+x^9)-(c^9+x^9)'(c^9-x^9)}{(c^9+x^9)^2} \\ f'(x)=\frac{(-9x^8)(c^9+x^9)-(c^9+x^9)'(c^9-x^9)}{(c^9+x^9)^2}\]

OpenStudy (anonymous):

n

OpenStudy (freckles):

so you still need to differentiate (c^9+x^9)

OpenStudy (freckles):

with respect to x of course

OpenStudy (anonymous):

yes i understand it now i just solved it and it was right thank you

OpenStudy (freckles):

ok cool

OpenStudy (freckles):

i said c isn't zero c is a constant and it could be 0 but they want the derivative formula for any constant c

OpenStudy (freckles):

and ok i know you get so i will shut up now

OpenStudy (anonymous):

lol its k thank you again